Mother gave Claire and Jane each a certain amount of money to spend. If Jane spent all her money and Claire spect 3 times as much, Claire would have $100 left. If Jane spent all her money and Claire spent 1/3 of what Jane spent, Claire would have $580 left. Find the total amount of money given to Claire and Jane.
Rmks: Pls try not to use algebra.

Answer by stanbon(48535)

Mother gave Claire and Jane each a certain amount of money to spend.
Clair got "C" ; Jane got "J".
Total received: C + J
----------------------------------
If Jane spent all her money and Claire spent 3 times as much,
Claire would have $100 left.
C - 3J = 100
----------------------------------
If Jane spent all her money and Claire spent 1/3 of what Jane spent, Claire would have $580 left.
C - (1/3)J = 580
-------------------------

Find the total amount of money given to Claire and Jane.
Claire had 3*Jane's amount + $100
Claire had (1/3)*Jane's amount + 580
----------------------------
Those amounts must be the same:
3J + 100 = (1/3)J + 580
(8/3)J = 480
(1/3)J = 60
J = $180 (amount Jane received from her mother)
Since Claire had 3*J + 100, C = 3*180+100 = $640
=========
Total received from mother was 580 + 180 = $760
=================
Cheers,
Stan H.